Service Department Charges and Activity Bases Middler Corporation, a manufacturer of electronics and communications systems, uses a service department charge system to charge proft centers with Computing and Communications Services (CCS) service department costs. The following table identifies an abbreviated list of service categories and activity bases used by the CCS department. The table also includes some assumed cost and activity base quantity information for each service for October. CCS Service Budgeted Activity Base Quantity Categony Activity Base Number of calls Number of devices monitored Number of user accounts Budgeted Cost Help desk $113,620 635,800 72,000 140,800 3,800 9,350 7,200 8,800 Network center Electronic mail Smartphone support Number of smartphones issued one of the profit centers for Middler Corporation is the Communication Systems (COMM) sector. Assume the following information for the COMM sector: The sector has 3,000 employees, of whom 40% are office employees. An the office employees have been issued a smartphone, and 80% of them have a computer on the network. 90 percent of the employees with a computer also have an e-mail account The average number of help desk calls for October was 1 call per individual with a computer. There are 260 additional printers, servers, and peripherals on the network beyond the personal computers a. Determine the service charge rate for the four CCS service categories for October. Round your answers to two decimal places Previous Next Check My Work
media%2Fb67%2Fb67fc757-0ad9-490d-bfd9-dc
Answer #1

hi, please find below the answer , let me know if you need any clarification -

Answer a)
i ii iii=i/ii
Category Activity base Budgeted cost Budgeted activity base quantity Service charge per unit
Help desk Number of call 113620 3800 $          29.90
Network center Number of device monitoring 635800 9350 $          68.00
Electronic mail Number of user account 72000 7200 $          10.00
Smartphone support number of smartphones issue 140800 8800 $          16.00
CCS service category Service charge rate
Help desk $                                             29.90 per call
Network center $                                             68.00 per divice monitoring
Electronic mail $                                             10.00 per user or e-mail account
Smartphone support $                                             16.00 per smartphones issued
Answer b)
Computation of charges for COMM sector
We have to first compute activity base for COMM sector
Given that-
Total number of employee 3000
Office employee =3000*40% 1200
Number of smart phone issued = 1200
Computer on network =1200*80% 960
Number of user account = 960*90% 864
Number of call = 960
Number of device monitored 260
October charges to COMM
Service charge rate Unit comsumed Cost
Help desk 29.9 960 28704
Network center 68 260 17680
Electronic mail 10 864 8640
Smartphone support 16 1200 19200
74224